# Quantitative and Predictive Folding Models from Limited Single-Molecule Data Using Simulation-Based Inference

### v1

@@ VIII Experimental Setup
The DNA hairpin data presented in this work were first presented in Ref.[46]. Briefly, the hairpin was attached to kilobase-length DNA handles containing either a biotin or digoxigenin molecule at their 5’-end using autosticky polymerase chain reaction (PCR). This construct was then attached to sub-micron diameter beads coated with either avdin or anti-digoxigenin, allowing the DNA hairpin to be held between two independently controlled optical traps. Tension could then be applied to the hairpin by increasing the distance between the two optical traps, which were separated until the hairpin began to spontaneously unfold. Once the hairpin was observed to occupy its folded and unfolded states with approximately equal probability, the distance between the two traps was fixed and measurements of the hairpin folding were collected by monitoring the separation between the two trapped beads. In order to measure the hairpin dynamics at constant force, we employed a passive force clamp by adjusting the power of one of the optical traps [47].
@@
IX Deconvolution
To deconvolve the point spread function from the hairpin data, we used a modified version of a deconvolution routine previously used on hairpin data [7]. The measured extension distribution of the hairpin $p(x)$ and the point spread function $S(x)$ determined from the reference construct were first convolved with a Gaussian function with a standard deviation of 2 $nm$ to smooth the data. We then made the initial guess of the deconvolved distributions $p^{(0)}_{\textrm{deconv}}(x)=p(x)$, and calculated their convolution with the point spread function $p^{(0)}_{\textrm{conv}}=p^{(0)}_{\textrm{deconv}}(x)\otimes S(x)$. We then computed the free energy landscape corresponding to the measured and convolved distributions as $G(x)=-log(p(x))$ and $G^{(0)}_{\textrm{conv}}(x)=-log(p^{(0)}_{\textrm{conv}}(x))$ respectively, subtracting the minimum free energy value from each landscape so that both were strictly positive. We then updated our estimate of the deconvolved distribution iteratively using the difference between the two free energy landscapes:
$p^{(i+1)}_{\textrm{deconv}}(x)=p^{(i)}_{\textrm{deconv}}(x)\cdot e^{-(G(x)-G^{(i)}_{\textrm{conv}}(x))}$
, renormalizing the updated guess of the deconvolved distribution after each iteration. We ran this process for 15 iterations, at which point the convolved and measured distributions matched well but the noise spikes common to deconvolution procedures had not yet appeared.

### v2

@@ IX Experimental Setup
The DNA hairpin data presented in this work were first presented in Ref. [30]. Briefly, the hairpin was attached to kilobase-length DNA handles containing either a biotin or digoxigenin molecule at their 5’-end using autosticky polymerase chain reaction (PCR). This construct was then attached to sub-micron diameter beads coated with either avdin or anti-digoxigenin, allowing the DNA hairpin to be held between two independently controlled optical traps. Tension could then be applied to the hairpin by increasing the distance between the two optical traps, which were separated until the hairpin began to spontaneously unfold. Once the hairpin was observed to occupy its folded and unfolded states with approximately equal probability (at a force of 13.9 pN), the distance between the two traps was fixed and measurements of the hairpin folding were collected by monitoring the separation between the two trapped beads. In order to measure the hairpin dynamics at constant force, we employed a passive force clamp by adjusting the power of one of the optical traps [21]. Data were sampled at 1 MHz and filtered at the Nyquist frequency.
@@
IX.2 Add riboswitch aptamer
We analyzed measurements of the add riboswitch aptamer first reported in Ref. [37]. Briefly, the aptamer was attached at each end to beads held in optical traps via kb-long duplex handles. Measurements at a constant force of 11.1 pN were done similarly to the hairpin, but with data sampled at 20 kHz. Our analysis used the first 5 s of the trajectory.
X Deconvolution
To deconvolve the point spread function from the hairpin data, we used a modified version of a deconvolution routine previously used on hairpin data [53]. The measured extension distribution of the hairpin $p(x)$ and the point spread function $S(x)$ determined from the reference construct were first convolved with a Gaussian function with a standard deviation of 2 $nm$ to smooth the data. We then made the initial guess of the deconvolved distributions $p^{(0)}_{\textrm{deconv}}(x)=p(x)$, and calculated their convolution with the point spread function $p^{(0)}_{\textrm{conv}}=p^{(0)}_{\textrm{deconv}}(x)\otimes S(x)$. We then computed the free energy landscape corresponding to the measured and convolved distributions as G(x)=log(p(x))G(x)=-log(p(x)) and $G^{(0)}_{\textrm{conv}}(x)=-log(p^{(0)}_{\textrm{conv}}(x))$ respectively, subtracting the minimum free energy value from each landscape so that both were strictly positive. We then updated our estimate of the deconvolved distribution iteratively using the difference between the two free energy landscapes: {align*} p^(i+1)_deconv(x) = p^(i)_deconv(x)⋅e^-(G(x)-G^(i)_conv(x)), renormalizing the updated guess of the deconvolved distribution after each iteration. We ran this process for 15 iterations, at which point the convolved and measured distributions matched well but the noise spikes common to deconvolution procedures had not yet appeared.